OpenStudy (loser66):

Convert \(P_2\) inti an inner product space by writing (x,y) =\(\int_0^1 x(t)\bar y(t) dt\) whenever x, y are in \(P_2\) and find a complete orthonormal set in that space Please, help

OpenStudy (loser66):

since x, y in P2, they have the same basis, right?

OpenStudy (loser66):

let say, the basis is {1,t,t^2}

OpenStudy (loser66):

unless convert it to dual basis (1,0,0), (0,1,0), (0,0,1) I don't know how to construct the orthonomal basis set

OpenStudy (anonymous):

i am probably way off base, but isn't \(\{1,t,t^2\}\) a complete orthonormal basis?

OpenStudy (anonymous):

the norm of each of them is 1 and it is certainly complete

OpenStudy (anonymous):

but i would not be $12 that it is correct, but it sure seems correct

OpenStudy (anonymous):

*bet

OpenStudy (loser66):

how can you know it is perpendicular to each other or not? they are l.i

OpenStudy (loser66):

As I said above, unless convert to standard basis, I have no way to prove those vectors orthogonal, then orthonomal set

OpenStudy (anonymous):

oh crap i guess you have to do that gramm schmidt business

OpenStudy (anonymous):

What are x and y?

OpenStudy (loser66):

polynomials in P2

OpenStudy (anonymous):

i take it they are polynomials of degree 2 but i could be wrong

OpenStudy (anonymous):

So \[ \sum_{n=0}^2a_nx^n \]

OpenStudy (anonymous):

But what would the conjugate be?

OpenStudy (loser66):

I think we proceed on complex space first and if argue on real then.

OpenStudy (anonymous):

So then the coefficients and \(t\) are complex numbers?

OpenStudy (loser66):

My prof said \(\bar y(t)\) means conjugate coefficients of y

OpenStudy (anonymous):

oh maybe it is not that bad put \(u_1=1\) then \(u_2=t-\int_0^11\times tdt=xt-1\)

OpenStudy (anonymous):

typo there i meant \(u_2=t-\int_0^11\times tdt=t-1\)

OpenStudy (anonymous):

you also have to divide by the norm or each vector, but the norm is one in each case so not an issue here

OpenStudy (anonymous):

no i am screwing this up

OpenStudy (anonymous):

What?\[ \int_0^1(1\times t)\;dt = \left[\frac{t^2}{2}\right]_0^1 = \frac 12 \]

OpenStudy (anonymous):

yeah i am messing this up totally but now what we have that, we have \(u_2=t-\frac{1}{2}\)

OpenStudy (anonymous):

Well, what I can say is we need \(\|x\| = (x_1,x_1) = 1^2=1\) and we need \((x_1,x_2)=0\).

OpenStudy (anonymous):

then you need the norm of \(t-\frac{1}{2}\) and divide by that

OpenStudy (anonymous):

Maybe try \(a\), \(bt\) and \(ct^2\), where \(a,b,c\in \mathbb C\) and try to solve for those coefficients?

OpenStudy (anonymous):

i was trying gramm schmidt are you supposed to do something else?

OpenStudy (loser66):

One question: What is Rieze Representation Theorem? does it work for this problem?

OpenStudy (anonymous):

i get if \(w_1=1\) then \(w_2=12t-6\)

OpenStudy (anonymous):

no idea but at least i have produced two that are orthogonal this gramm schmidt actually seems to work

OpenStudy (anonymous):

I am getting \((w_2,w_2) = 12\)... or am I suppose to do something weird to get \(\overline{w_2}\)?

OpenStudy (anonymous):

i am still messing up let me go slower

OpenStudy (anonymous):

I think I know what you're trying to do now.

OpenStudy (anonymous):

You are getting a projection using the provided inner product. But how would you ever get any complex coefficients doing this? Or does not not matter?

OpenStudy (anonymous):

i forgot to take the damned square root!

OpenStudy (anonymous):

ok finally i think i have the first two \(1\) and \(2\sqrt3(t-\frac{1}{2})\)

OpenStudy (anonymous):

i will probably make four or five mistakes before i find the third one

OpenStudy (anonymous):

i can walk you through what i did if you like but if you are supposed to use something other than gramm schmidt i don't know it

OpenStudy (loser66):

I understand what you are doing, my question is you are trying to find an orthonomal set on the given basis, how it relates to the (x,y) = int.....? how the dual basis y which supposed to be orthonomal dual of x is defined?

OpenStudy (loser66):

Using Gramm Schmidt to orthonomal the given basis is not hard, (just be patient), how to put it in use is what I want to know. :)

OpenStudy (anonymous):

maybe i don't understand the question i thought you were looking for an orthonormal basis

OpenStudy (anonymous):

You can use orthonormal basis to project one vector subspace onto another vector subspace. Projections are good for finding the solution with the least error, when the exact solutions are in another subspace.

OpenStudy (loser66):

I have a note from my prof, however, I don't understand what he was talking about. Let me copy it V = P_n define (x,y)= \(\int_a^b x(t)\bar y(t) dt\) an inner product on P_n Let f in P'_n given by [x,f]= \(x(t_0)\) in Complex By Riesz Representation Theorem, there exists unique P in P_n such that \(\int_a^b x(t)\bar P(t) dt = x(t_0)\) for every x in P_n P depends on n, a, b, \(t_0\) end note.